OpenStudy (anonymous):

Use the properties of exponents to simplify the expression. Write with positive exponents. Assume that all variables represent positive real numbers. x^(-2/5) times x^(7/5)

jhonyy9 (jhonyy9):

x^(-2/5 +7/5) =x^((7-2)/5) =x^(5/5) =x^1 =x